OpenStudy (blyeeeeee):

Find the angle measure that makes the statement true. Sin 25° = Cos ____ °

OpenStudy (sshayer):

\[\cos \left( 90-\theta) \right)=\sin \theta\]

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Look at the figure above. What is \(\sin A\)? \(\sin A = \dfrac{opp}{hyp}\) \(\sin A = \dfrac{y}{h} \) Now again using the figure above, what is \(\cos B\)? \(\cos B = \dfrac{adj}{hyp} \) \(\cos B = \dfrac{y}{h}\) In a right triangle with acute angles A and B, we have \(\sin A = \cos B\) Now notice what is the relation between the acute angles of a right triangle. The measures of the three angles of any triangle add up to 180 degrees. Since a right triangle has measure 90 degrees, that means the two acute angles of a right triangle have measures that add up to 90 degrees. That means that A + B = 90, and B = 90 - A. From this we derive an important relation in trigonometry. This is known as a co-function identity. \(\sin A = \cos (90^\circ - A)\) \(\cos A = \sin (90^\circ - A)\) This applies not only to sine and cosine, but to the other trig co-functions: tangent and cotangent secant and cosecant